# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ Abstract
The development of fault-tolerant quantum computers (FTQCs) is receiving increasing attention within the quantum computing community. Like conventional digital computers, FTQCs, which utilize error correction and millions of physical qubits, have the potential to address some of humanity’s grand challenges. However, accurate estimates of the tangible scale of future FTQCs, based on transparent assumptions, are uncommon. How many physical qubits are necessary to solve a practical problem intractable for classical hardware? What costs arise from distributing quantum computation across multiple machines? This paper presents an architectural model of a potential FTQC based on superconducting qubits, divided into discrete modules and interconnected via coherent links. We employ a resource estimation framework and software tool to assess the physical resources required to execute specific quantum algorithms compiled into their graph-state form and arranged onto a modular superconducting hardware architecture. Our tool can predict the size, power consumption, and execution time of these algorithms as they approach the utility scale based on explicit assumptions about the physical layout, thermal load, and modular connectivity of the system. Using this tool, we assess the total resources in the proposed modular architecture and highlight the impact of trade-offs, intermodule connectivity, latency, and space-time resource requirements.
Rapid advancement in the development of FTQCs heralds a new era in which universal machines with millions of qubits can surpass classical computers in addressing specific problems (see [1, 2, 3, 4, 5, 6] for recent overviews and developments). At the same time, the size, connectivity, and fidelity of today’s noisy qubit devices are continuously improving [7, 8, 9]. These platforms are already considered the foundational elements of early FTQCs. They employ redundancy to suppress noise using a technique known as quantum error correction (QEC) [10, 11, 12, 4, 13, 9].
@@
The complexity and asymptotic scaling of specific fault-tolerant algorithms, including Shor’s factoring [2, 6], quantum signal processing [14] (QSP), and ground state energy estimation [15] (GSEE), are well understood. However, research on the magnitude and scaling of the physical resources needed to execute such computations is limited. Furthermore, the overhead and space-time resource requirements for physical hardware are even less well understood and depend on various choices related to platforms and micro-architectures. Scaling tangible resources such as energy, time, and space based on specific assumptions about a potential machine is crucial for evaluating the feasibility of methods to develop FTQC.
Here, we present a transparent fault-tolerant architectural model based on square lattices of superconducting qubits utilizing the surface code [10, 12, 16, 17, 18], allocated to the essential elements of FTQC: logical Clifford gates through lattice surgery, magic or $T$ state distillation (MSD) [11, 12, 19] for non-Clifford gates, as well as considering the queuing and routing of specific resources. We also establish explicit assumptions regarding the size of a single module and the penalties associated with executing a fault-tolerant algorithm that exceeds the capacity of a single module to quantify the impact of coherently distributing a quantum computation.
@@
Resource estimation tools exist based on the “T-counting” method [12, 20, 21, 22, 23]. These footprint-style estimation tools synthesize the logical circuit into Clifford+$T$ gates (or other universal gate sets), rescale the space-time volume according to the desired level of parallelization, and derive physical space-time quantities based on the $T$-count and circuit depth, along with assumptions regarding $T$-distillation and hardware timescales. This methodology lacks targeted estimates specific to Quantum Processing Unit (QPU) layouts, provides insufficient estimates for surface code distance, and may overestimate the upper limits of space-time resources.
@@
We also present application-specific resource estimates for the proposed fault-tolerant superconducting qubit architecture. These estimates are based on a graph state method detailed in section I.1. This framework aims to assist in designing and assessing the feasibility of next-generation FTQCs while quantifying the trade-offs associated with hardware implementations. We also discuss an accompanying open-source software called Rigetti Resource Estimations [24] (RRE), which we used to generate all the results presented in section III for a selection of test algorithms outlined in section II. We will explore potential future work in section IV and provide a conclusion in section V.
@@
I A primer for algorithm execution, architecture designs, and resource estimation on FTQCs
@@
The execution of algorithms on an FTQC fundamentally differs from merely executing the gates of a quantum circuit in the order they appear. By definition, FTQCs depend on the underlying QEC [10, 11, 12, 4, 13, 9] scheme they utilize. FTQCs employ noisy physical qubits, encoded within an error correction scheme, to create logical qubits with exponentially suppressed error rates.
@@
The error-correcting scheme we consider here is the surface code [10, 12, 16, 17], which is implemented on square arrays of physical qubits. In an array, qubits are categorized into physical data or ancilla qubits, with each connected data qubit coupled to four ancilla qubits and vice versa. Ancilla qubits facilitate repeated composite parity measurements on their four neighboring data qubits of form $X_{1}X_{2}X_{3}X_{4}$ or $Z_{1}Z_{2}Z_{3}Z_{4}$. The composite parity measurements on the data qubits can be performed by entangling them with the ancilla through CNOTs and measuring the ancilla in either the Pauli $X$ or $Z$ basis. Physical errors on qubits in these arrays result in patterns of bit flips among the parity measurements known as syndromes. Tracking the outcomes of the parity measurements, the resulting error syndromes, and mapping them to the most probable error event is known as decoding. In practical terms, the corrections applied after decoding can be accomplished through judicious syndrome tracking and implementing corrections in software after running the algorithm (potentially incurring time delays due to decoder latency).
@@
The price we pay for this encoding is that the resulting logical qubits can only perform a limited set of operations, fault-tolerantly, specifically the Clifford set. At least one non-Clifford operation must be engineered in the FTQC to maintain universality and execute arbitrary fault-tolerant algorithms. Typically, and for this paper, we consider the $T$ gate: a $\pi/4~{}R_{z}$-rotation. However, forcing an error-correcting code to produce an operation outside its native subspace, like a $T$-gate, involves significant overhead, necessitating the dedication of substantial portions of the FTQC to MSD or, more recently, cultivation [11, 12, 19].
@@
Figure 1: (top) Fault-tolerant algorithm execution involves segmenting the application into logical Clifford and non-Clifford portions. Clifford portions are accomplished through a set of measurements of logical qubits facilitated by an ancilla bus. Due to the probabilistic nature of measurement outcomes, classical feedback and feedforward are incorporated between the schedules of these operations. Non-Clifford portions are accomplished using magic state distillation. (bottom) This results in a layout of the logical micro-architecture shown for illustration purposes only. Each square represents a surface code patch of distance $d$. A register of logical memory qubits (blue tiles) connects through a bus of ancilla qubits (yellow tiles). Below it, a separate bus is designated for queuing $T$-states created in an array of $T$-factories. This architecture can be distributed across multiple machines, provided they are connected via the ancilla bus and each module consumes local $T$-states. We implement a modified version with a wrapped, more practical quantum register, as seen in fig. 6.
@@
Due to the restrictions imposed by the error-correcting code, we ultimately divide the algorithm into operations consisting of Clifford+$T$ gates executed sequentially. This separation of functional operations leads to a specialization of the fundamental architecture: Certain sections of the FTQC are dedicated solely to performing the original logical Clifford operations through multi-product (qubit) Pauli measurements—a process called lattice surgery [25, 26, 27, 28, 12]. For example, the multi-product Pauli measurements can be understood as executing $n$-body $CZ$ and single-qubit Clifford gates on the quantum state, leading to a formalism we later explore. Additionally, there are designated areas for generating and queuing $T$ states produced via MSD. An example of this layout is shown schematically in fig. 1.
@@
We adapt this logical quantum operation model to a fixed yet extensible superconducting qubit architecture, modified from fig. 1 and consisting of a physical layout and square-lattice connectivity. We estimate the execution of selected algorithms on this architecture. To facilitate this, we generate schedules of quantum operations using a graph-state-based approach, as proposed in several studies [29, 30, 31, 32, 33, 34, 35, 36, 37] and further refined below. This approach yields several abstractions that aid in the distribution of the algorithm across the fixed hardware. Namely, we can manipulate a graph or subgraphs to customize the entire algorithm execution on an FTQC composed of distinct yet interconnected modules.
@@
One can implement the original quantum algorithm using the algorithm-specific graph (ASG) formalism [38, 31, 32, 39] and associated schedules of operations [34, 33], which apply only logical measurements and classical feedforward. The ASG formalism is a cost-efficient variant of the Measurement-Based Quantum Computing (MBQC) framework — the latter was introduced in [35, 40]. MBQC is particularly appealing for distributed FTQC because it requires only an algorithm-specific resource, such as the graph states and measurements on logical qubits. We will explore and utilize a specific form of MBQC as outlined in [31] for compilation purposes. We prepare the graph-state resource by using lattice surgery to perform entangling operations between qubits on a register of memory qubits via a single bus of ancillary qubits. Due to the probabilistic nature of quantum measurements, we must track the set of measurement results classically using a process called Pauli frame tracking [41, 42, 43, 33]. Next, we teleport local $T$-states to the register of memory qubits to apply the non-Clifford portion in conjunction with the application of the Pauli Frame corrections of the algorithm. Finally, we apply any corrections flagged by syndromes tracked by the decoder, resulting from physical errors in the hardware. We present a mathematical description of the ASG formalism in section I.1.
@@
How can we arrange the elements of the graph-state-based execution of the algorithm across physical components at scale? In large-scale classical computations, there are limits to the amount of computation that can be executed on a single processor within a specific time frame. These limits arise from the processing power that can be integrated into an individual chip, ultimately influenced by physical constraints and manufacturing yields. Consequently, computations are often distributed across multiple processors and machines. This distribution incurs penalties, such as higher latency or reduced bandwidth, but it enables us to leverage increased overall computational power. We can draw an analogy to distributed FTQC, which involves a network of interconnected compute modules. Manufacturing yields limit the quantum computing power of each module. These operational (thermal) and manufacturing constraints restrict the number of qubits that can be incorporated into a single QPU housed within a refrigeration module. This prompted us to propose a logical macro-architecture composed of several modules described in section I.2.
@@
In contrast to the flexible connectivity and relatively slow speeds found between compute modules, QPUs offer fixed native connectivity and faster gate speeds. In section I.3, we describe hardware micro-architecture, how it can be customized to accommodate specific hardware constraints, and a physical rationale for why the number of qubits in a single compute module should be around one million. We propose sparsely interconnecting the modules, which may result in reduced fidelity, lower bandwidth, or slower entangling gate speeds. We will refer to the connection between modules with the surface code distance of $d$ as a single "pipe." We will abstract the transduction efficiency as an interaction time, characterized by the "intermodule tock," and quantify how this time impacts the algorithm’s overall runtime (we assume that the pipes share identical noise characteristics and code distance with the rest of FTQC).
@@
Finally, we present our compilation and estimation methodology based on user inputs in RRE in section I.4.6. Further details regarding the output glossary of RRE, our circuit decomposition strategy, and the verification of RRE outputs can be found separately in section A.1, section A.2.4, and section A.3, respectively.
I.1 Mathematical description of the graph-state formalism
A mathematical graph, denoted as $\mathcal{G}$, can be defined by its vertices and edges. The graph state associated with a quantum (Clifford) state is constructed by assigning each vertex to the qubits initialized to $\ket{+}$ and each edge to an entangling $CZ$ operation that links the qubits on the corresponding vertices. We often refer to the qubit corresponding to a vertex as a "node" and interchange the terms "graph" and "graph state" for brevity. The nodes will undergo non-Pauli basis measurements corresponding to the original circuit’s non-Clifford gates and some $X$-basis measurements to facilitate the teleportation of unknown inputs into the graph state [31].
@@ I.1 Mathematical description of the graph-state formalism
The ASG formalism provides measurement schedules for two main stages to fully execute the fault-tolerant algorithm: one schedule details the preparation of the graph through lattice surgery (corresponding to the original nontrivial Cliffords, first stage, in fig. 1(top)) and another that outlines the necessary single-qubit measurements in non-Clifford bases (corresponding to the original non-Cliffords, second stage, in fig. 1(top)). As discussed previously, we also need to track the Pauli frame corrections in the software. We denote the nested set of stabilizer measurement schedules that "prepare" the graph in the first stage as $S^{\text{prep}}_{\mathcal{G}}$. Conversely, we denote the nested set of non-Clifford measurement schedules that "consume" the graph state in the second stage as $S^{\text{consump}}_{\mathcal{G}}$. Both schedules consist of subsets (subschedules) that specify the order of measurements, with each subset containing the nodes designated for simultaneous measurements [44, 45, 31]. Furthermore, the essential stages of MSD, routing $T$-states, and decoding can occur or be postponed at either of the two stages, depending on the requirements.
@@
At a high level, the ASG formalism requires ingesting the fault-tolerant algorithm as single- and two-qubit gates. We employ a straightforward method for decomposition, compilation, and execution of the algorithm, segmenting it into equal-width, recurring subcircuits along the time axis. These subcircuits are referred to as "widgets," have their own (sub)graphs, and reconstruct the original algorithm when "stitched" together (see also section I.4.6). We refer to this decomposition process as "widgetization," detailed in section A.2.4. We will examine the graph properties of these widgets and the scheduling process for preparing and consuming them on the proposed hardware architecture.
@@
I.2 High-level macro-architecture for modular graph state processing
This section describes a macro-architectural model utilizing superconducting qubits, enabling us to estimate the resources necessary for executing the fault-tolerant algorithm while understanding the associated trade-offs.

### v3

# Superconducting qubits in the millions: the potential and limitations of modularity

@@ Abstract
The development of fault-tolerant quantum computers (FTQCs) is receiving increasing attention within the quantum computing community. Like conventional digital computers, FTQCs, which utilize error correction and millions of physical qubits, have the potential to address some of humanity’s grand challenges. However, accurate estimates of the tangible scale of future FTQCs, based on transparent assumptions, are uncommon. How many physical qubits are necessary to solve a practical problem intractable for classical hardware? What costs arise from distributing quantum computation across multiple machines? This paper presents an architectural model of a potential FTQC based on superconducting qubits, divided into discrete modules and interconnected via coherent links. We employ a resource estimation framework and software tool to assess the physical resources required to execute specific quantum algorithms compiled into their graph-state form and arranged onto a modular superconducting hardware architecture. Our tool can predict the size, power consumption, and execution time of these algorithms based on explicit assumptions about the system’s physical layout, thermal load, and modular connectivity. We assess the resources needed for quantum computation examples that serve as building blocks of proposed applications, quantifying the architectural bottlenecks and trade-offs that remain to be addressed to deliver utility.
I Introduction
Recent studies suggest that future idealized FTQCs with 1-10 million physical qubits can address classically intractable problems in integer factoring [25] and molecular chemistry [36, 26], completing their assigned tasks in a matter of days. At the same time, the size, connectivity, and fidelity of today’s noisy physical qubit devices are continuously improving, and now support some of the fundamental building blocks of fault tolerance [14, 17, 1]. These emerging platforms employ quantum error correction (QEC) codes [20, 10, 40, 39, 88, 2, 84, 1] to reduce logical errors at the cost of additional resources: more space, meaning physical qubits, and more time needed to operate the code.
@@ I Introduction
The complexity of fault-tolerant (FT) algorithms, including Shor’s factoring [82], quantum signal processing [43] (QSP), and ground state energy estimation [37], is well understood. However, research on the magnitude and scaling of the physical resources required to execute such computations is limited. Furthermore, the overhead and space-time resource requirements for physical hardware are even less well understood and depend on platform and micro-architecture choices. Scaling tangible resources such as energy, space, and time, based on specific assumptions about a potential machine, is crucial for evaluating the feasibility of methods for developing FTQC.
@@
Fault-tolerant resource estimation tools currently exist, but are restricted to using the $T$-counting” method [39, 61, 56, 7, 53] only. These footprint-style estimation tools synthesize the logical circuit into Clifford+$T$ gates (or other universal gate sets), rescale the space-time volume according to the desired level of parallelization, and derive physical space-time quantities based on the $T$-count and circuit depth, along with assumptions regarding $T$-distillation and hardware timescales. This methodology lacks targeted estimates specific to Quantum Processing Unit (QPU) layouts, provides insufficient estimates for surface code distance, and may overestimate the upper limits of space-time resources.
Here, we present a transparent, FT architectural model for square-lattice superconducting-qubit systems that uses the surface code [20, 39, 52, 18, 58]. This model supports the essential elements of FTQC: logical Clifford gates via lattice-surgery-based approaches [30, 9, 38, 19, 39], magic- or $T$-state distillation (MSD) [40, 39, 21] for non-Clifford gates, and queuing and routing of specific resources. We also establish explicit assumptions about the size of a single module and the penalties for executing an FT algorithm that exceeds its capacity, allowing us to quantify the impact of coherently distributing a quantum computation.
@@
We also present application-specific resource estimates for the proposed FT superconducting qubit architecture. These estimates are derived from a graph-state method detailed in section˜II.1. This framework aims to assist the design and feasibility assessment of next-generation FTQCs while quantifying the trade-offs associated with hardware implementations. We also discuss an accompanying open-source software called Rigetti Resource Estimations [77] (RRE), which we used to generate all the results presented in section˜IV for a selection of test algorithms outlined in section˜III. We will explore potential future work in section˜V and provide a conclusion in section˜VI.
@@
II A primer for algorithm execution, architecture designs, and resource estimation on FTQCs
@@
The execution of algorithms on an FTQC fundamentally differs from merely executing the gates of a quantum circuit in the order they appear. By definition, FTQCs depend on the underlying QEC scheme they use - see, for example, [20, 40, 39, 2, 84, 1]. FTQCs employ noisy physical qubits, encoded within an error-correction scheme, to create logical qubits with exponentially suppressed error rates.
@@
The error-correcting scheme we consider here is the surface code [20, 39, 52, 18], implemented on square arrays of physical qubits. In an array, physical qubits are categorized as data or ancilla qubits, with each ancilla qubit coupled to four data qubits and vice versa. Physical ancilla qubits enable repeated composite parity measurements on their four neighboring data qubits in the form $X_{1}X_{2}X_{3}X_{4}$ or $Z_{1}Z_{2}Z_{3}Z_{4}$. These composite parity measurements are performed by entangling the physical data qubits with the ancilla via CNOTs and measuring the ancilla in either the Pauli $X$ or $Z$ basis. Physical errors on the physical qubits in these arrays result in patterns of bit flips among the parity measurements, known as syndromes. Tracking the outcomes of the parity measurements and the resulting error syndromes, and mapping them to the most probable error event, is known as decoding. In practice, corrections can be applied after decoding through judicious syndrome tracking and software corrections after running the algorithm (potentially incurring time delays due to decoder latency).
@@
One price we pay for this encoding is that the resulting logical qubits can perform only a limited set of FT operations, specifically the Clifford set. At least one non-Clifford operation must be engineered into the FTQC to achieve universality and execute arbitrary algorithms fault-tolerantly. We consider the $T$ gates (a $\pi/4~R_{z}$ rotation) as the non-Clifford operations choice in this paper. However, forcing an error-correcting code to produce an operation outside its native subspace (such as a $T$ gate) involves significant overhead, necessitating the dedication of substantial portions of the FTQC to MSD or, more recently, cultivation [40, 39, 21] to produce $T$ gates.
@@
Figure 1: (top) Fault-tolerant algorithm execution involves segmenting the application into logical Clifford and non-Clifford portions. Clifford operations are performed using a set of measurements on logical memory qubits, facilitated by auxiliary buses. Because measurement outcomes are probabilistic, classical feedback and feedforward are incorporated into operation schedules. Non-Clifford portions are implemented using $T$-states distilled by $T$-factories. (bottom) This results in our proposed logical micro-architecture, detailed in section˜II.3. The architecture includes graph preparation, consumption, local $T$-factories, $T$-state and Bell-state queuing, with some parallelization. We arrange logical qubits on a square lattice of rotated surface code patches of distance-$d$ made of physical qubits. A left portion of the module is allocated for the logical memory (blue tiles). It also contains an “auxiliary bus” of logical auxiliary qubits needed for graph preparation, consumption, $T$, and Bell-state routing. Yellow tiles represent the auxiliary bus, followed by lighter-yellow “auxiliary++ bus” required for routing. To its left, a separate bus is designated for continuous queuing of $T$-states distilled in an array of $T$-factories on the opposite side of the module (teal tiles) — see also Parameters 6 and 8 of appendix˜A. This architecture can be distributed across multiple modules, provided they are coherently connected via the auxiliary bus and each module consumes local $T$-states.
@@
Due to the constraints imposed by the error-correcting code, we ultimately decompose the algorithm into a sequence of Clifford+$T$ gates. This separation of functional operations leads to a specialization of the fundamental architecture: certain sections of the FTQC are dedicated solely to performing the original logical Clifford operations through multi-product (-qubit) Pauli measurement operations or MPPOs — a process based on the lattice surgery [30, 9, 38, 19, 39]. For example, MPPOs can be understood as applying $n$-body CZCZ gates and single-qubit Clifford gates to the quantum state, thereby leading to a formalism that we later explore. Additionally, there are designated areas for generating and queuing $T$ states produced via MSD. Our proposed logical layout for this operational model is presented schematically in fig.˜1.
@@
We adapt this logical-operation model and the fixed yet extensible micro-architecture of fig.˜1 to a physical layout of superconducting qubits with square-lattice connectivity. We estimate the execution of selected algorithms on this architecture. To this end, we generate schedules of quantum operations using a graph-state-based approach, as proposed in [28, 4, 87, 76, 42, 68, 69, 70] and further refined below. This approach yields several abstractions that facilitate distributing the algorithm across fixed hardware. Specifically, we can manipulate graphs or subgraphs to tailor the overall algorithm’s execution within an FTQC composed of distinct yet interconnected modules.
@@
One can implement the original quantum algorithm using the algorithm-specific graph (ASG) formalism [63, 87, 8] and associated schedules of operations [42, 76], which apply only logical measurements and classical feedforward. The ASG formalism is a resource-efficient variant of the Measurement-Based Quantum Computing (MBQC) framework [68, 57]. MBQC is particularly appealing for distributed FTQC because it requires only an algorithm-specific resource, such as the graph states and measurements on logical qubits. We will explore and utilize a specific form of MBQC as outlined in [87] for compilation purposes. We prepare the graph-state resource using lattice-surgery-based processes to entangle logical qubits in a memory register, facilitated by a single bus of logical auxiliary qubits. Due to the probabilistic nature of quantum measurements, we must track the set of measurement results classically using a process called Pauli frame tracking [33, 60, 71, 76]. Next, we move locally-distilled $T$-states, through patch deformations[39, 30], to the register of logical memory qubits to apply the non-Clifford portion in conjunction with the Pauli Frame corrections of the algorithm. Finally, we apply any corrections flagged by syndromes tracked by the decoder, resulting from physical errors in the hardware. We present a mathematical description of the ASG formalism in section˜II.1.
@@
We need to arrange the elements of the graph-state-based algorithm execution across physical components at scale. In large-scale classical computations, there are limits to the amount of computation that can be performed on a single processor within a given time frame. These limits stem from the processing power that can be integrated into a single chip, ultimately constrained by physical limits and manufacturing yields. Consequently, computations are often distributed across multiple processors and machines. This distribution incurs penalties, such as higher latency or reduced bandwidth, but it enables greater overall computational power. We can draw an analogy to distributed FTQC, which involves a network of interconnected compute modules. Manufacturing yields limit the quantum computing power of each module. These operational (thermal) and manufacturing constraints restrict the number of physical qubits that can be incorporated into a single QPU housed within a refrigeration module. This prompted us to propose a logical macro-architecture composed of several modules, as described in section˜II.2.
@@
In contrast to the flexible connectivity and relatively slow speeds between compute modules, QPUs based on superconducting qubits offer fixed connectivity and faster gate speeds. In section˜II.3, we describe the hardware micro-architecture, explain how it can be customized to accommodate specific hardware constraints, and provide a physical rationale for fixing the number of physical qubits in a single compute module to be around one million. We propose sparsely and coherently interconnecting the modules, which may result in reduced fidelity, lower bandwidth, or slower entangling gate speeds. We will refer to the “coherent interconnect” between modules with the surface code distance of $d$ as a single “pipe.” We will abstract the transduction efficiency as an interaction time, characterized by the “intermodule tock,” and quantify how this time impacts the algorithm’s overall runtime. We assume that the pipes share identical noise characteristics and code distance with the rest of FTQC.
@@
We present our compilation and estimation methodology for RRE, based on user inputs, in section˜II.4.7. Further details on the RRE output glossary, our circuit decomposition strategy, and the verification of RRE outputs are provided separately in appendix˜A, section˜B.0.4, and appendix˜C, respectively.
Setting aside the technical details of our proposed logical layouts, it is important to note that the fixed intra-module micro-architecture in fig.˜1 and the inter-module macro-architecture in section˜II.2 were not tailored to any specific application. Instead, we adapted the overall architecture to the strengths and limitations of our superconducting hardware and the graph-state-based algorithm execution. We will then evaluate its performance on selected applications. In fig.˜1, the bilinear pattern in the memory and auxiliary buses arises from the ease of scheduling the full FT algorithm using graph states. $T$-factories are kept local to modules and connected via another bus primarily because practical algorithms require a large number of $T$ operations, and coherent interconnects are slow for $T$-teleportations. However, we assume that coherent interconnects are effective for less frequent Bell-state-based teleportation of the graph state. This is why we propose the distributed macro-architecture described in section˜II.2; the graph state would not fit into a single module for practical algorithms. Future design improvements could include alternative memory and bus structures, such as city-style patch layouts and more efficient distinct QEC codes for memories, to reduce reliance on many coherent interconnects, which may be more suitable for certain applications, as suggested by our results in section˜IV.
II.1 Mathematical description of the graph-state formalism
[…]
@@ II.1 Mathematical description of the graph-state formalism
The ASG formalism provides measurement schedules for two main stages to fully execute the FT algorithm: one schedule details the preparation of the graph through lattice surgery (corresponding to the original nontrivial Cliffords, first stage, in fig.˜1(top)) and another outlines the necessary single-qubit measurements in non-Clifford bases (corresponding to the original non-Cliffords, second stage, in fig.˜1(top)). We also need to track Pauli frame corrections in the software. We denote the nested set of stabilizer measurement schedules that “prepare” the graph in the first stage as $S^{\text{prep}}_{\mathcal{G}}$ and refer to it as the preparation schedule for the graph state. $S^{\text{prep}}_{\mathcal{G}}$ specifies how to initialize $\mathcal{G}$, the graph state, based on the original entangling operations and other step-by-step requirements [42, 41]. Conversely, we denote the nested set of non-Clifford measurement schedules that “consume” the graph state in the second stage as $S^{\text{consump}}_{\mathcal{G}}$ and refer to it as the consumption schedule. $S^{\text{consump}}_{\mathcal{G}}$ specifies how to perform logical $R_{Z}(\theta)$-basis measurements (for arbitrary non-Clifford angle $\theta$) relevant to executing the original non-Cliffords [76, 86, 72]. Both schedules consist of subsets or subschedules that specify the order of measurements, with each subset containing the nodes designated for simultaneous measurements [41, 86, 87]. Furthermore, the essential stages of MSD, routing $T$-states, and decoding can occur at either stage or be postponed, depending on the requirements.
@@
At a high level, the ASG formalism requires the FT algorithm to be ingested as single- and two-qubit gates. We employ a straightforward method for decomposing, compiling, and executing the algorithm, segmenting it into equal-width, recurring subcircuits along the time axis. These subcircuits are called “widgets,” have their own (sub)graphs, and reconstruct the original algorithm when “stitched” together (see also section˜II.4.7). We refer to this decomposition process as “widgetization,” detailed in section˜B.0.4. We will examine the graph properties of these widgets and the scheduling process for preparing and consuming them on the proposed hardware architecture.
@@
II.2 High-level macro-architecture for modular graph state processing